OpenStudy (anonymous):

How do you solve If the sum of three consecutive even integer is 60 what is the first of the three even integers

OpenStudy (anonymous):

\[n, n+2, n+4\] addition results to 60 \[\left( n \right)+\left( n+2 \right)+\left( n+4 \right)=60\] \[3n + 6 = 60\] \[3n =54 \] \[n= \frac{ 54 }{ 3 }\]

OpenStudy (anonymous):

\[n = 18\] the numbers are 18, 20, 22